Presented here are summarized data from the balance sheets and income statements of Wiper, Inc.: WIPER, INC. Condensed Balance Sheets December 31, 2017, 2016, 2015 (in millions) 2017 2016 2015 Current assets $ 793 $ 1,021 $ 883 Other assets 2,428 1,935 1,734 Total assets $ 3,221 $ 2,956 $ 2,617 Current liabilities $ 592 $ 845 $ 742 Long-term liabilities 1,607 1,070 942 Stockholders’ equity 1,022 1,041 933 Total liabilities and stockholders' equity $ 3,221 $ 2,956 $ 2,617 WIPER, INC Selected Income Statement and Other Data For the year Ended December 31, 2017 and 2016 (in millions) 2017 2016 Income statement data: Sales $ 3,065 $ 2,928 Operating income 311 325 Interest expense 99 80 Net income 236 231 Other data: Average number of common shares outstanding 42.8 48.2 Total dividends paid $ 65.0 $ 53.8

e. If Wiper's stock had a price/earnings ratio of 12 at the end of 2017, what was the market price of the stock? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Hint: 5.51 is incorrect

f. Calculate the cash dividend per share for 2017 and the dividend yield based on the market price calculated in part e. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Cash dividends per share: $1.52

Dividend Yield: ?

h. Assume that accounts receivable at December 31, 2017, totaled $324 million. Calculate the number of days' sales in receivables at that date. (Use 365 days a year. Do not round intermediate calculations.)

i. Calculate Wiper's debt ratio and debt/equity ratio at December 31, 2017 and 2016. (Round "Debt ratio" to 1 decimal place and "Debt/equity ratio" to the nearest whole percent.)

2017 2016
Debt Ratio (0.7 is Incorrect) (0.6 is Incorrect)
Debt / Equity Ratio 215% Correct (65% is Incorrect)
Answer #1
[$ in millions] 2017 2016 2015
Current assets $             793 $        1,021 $           883
Other assets $         2,428 $        1,935 $        1,734
Total assets $         3,221 $        2,956 $        2,617
Current liabilities $             592 $            845 $           742
Long term liabilities $         1,607 $        1,070 $           942
Stockholders' equity $         1,022 $        1,041 $           933
Total liabilities and stockholders' equity $         3,221 $        2,956 $        2,617
Sales $         3,065 $        2,928
Operating income $             311 $            325
Interest expense $               99 $              80
Net income $             236 $            231
Average number of common shares $               43 $              48
Total dividends paid $               65 $              54
ANSWERS:
e] EPS [NI/Number of shares] $           5.51 $           4.79
P/E ratio 12
Market Price = EPS * P/E = $         66.17
f] Cash dividend per share $           1.52
Dividend yield = 1.52/66.17 = 2.30%
h] Days' sales in receivables = 324*365/3065 = 38.58 Days
i] Debt ratio [Total debt/Total assets] 68.3% 64.8%
Debt/Equity ratio 215.2% 184.0%